OpenStudy (anonymous):

Find the inverse of the function y=e^x+3

OpenStudy (experimentx):

x = e^y+3 x - 3 = e^y ln(x - 3) = ln(e^y) y = ln(x - 3)

OpenStudy (anonymous):

thank you! That helps, I was close :)